### v2

# GuideGen: A Text-Guided Framework for Full-torso Anatomy and CT Volume Generation

Abstract
The recently emerging conditional diffusion models seem promising for mitigating the labor and expenses in building large 3D medical imaging datasets. However, previous studies on 3D CT generation have yet to fully capitalize on semantic and textual conditions, and they have primarily focused on specific organs characterized by a local structure and fixed contrast. In this work, we present GuideGen, a controllable framework that generates anatomical masks and corresponding CT volumes for the entire torso—from chest to pelvis—based on free-form text prompts. Our approach includes three core components: a text-conditional semantic synthesizer for creating realistic full-torso anatomies; a contrast-aware autoencoder for detailed, high-fidelity feature extraction across varying contrast levels; and a latent feature generator that ensures alignment between CT images, anatomical semantics and input prompts. To train and evaluate GuideGen, we compile a multi-modality cancer imaging dataset with paired CT and clinical descriptions from 12 public TCIA datasets and one private real-world dataset. Comprehensive evaluations across generation quality, cross-modality alignment, and data usability on multi-organ and tumor segmentation tasks demonstrate GuideGen’s superiority over existing CT generation methods.
1 Introduction
Figure 1: Overview of GuideGen. Given a medical prompt, GuideGen is able to synthesize anatomical structures that accurately reflect demographic or clinical details described in the prompt in three stages. In Stage-I, GuideGen synthesizes discrete semantic volumes that conforms to spatial features designated in the medical prompt; In Stage-II, GuideGen deploys a contrast-aware autoencoding module to accommodate large CT volumes while extracting detailed texture from all contrast levels; In Stage-III, GuideGen combines information of semantic latents derived from anatomies in Stage-I, image latents extracted in Stage-II and textual latents from the input medical prompt.
The acquisition of a large quantity of medical images and their corresponding labels has always been critical for modern medical image analysis. However, due to data privacy issues and laborious annotation work, such datasets are often unavailable publicly, undermining the performance of subsequent tasks [14]. Recently trending conditional generative methods offer a promising solution to these problems: By eliminating privacy concerns and human interventions, they can generate vivid samples to an arbitrary scale, which provide significant boost to the performance of downstream applications like image segmentation and classification [31, 54, 29, 25].
Current conditional generative approaches can be categorized based on the nature of the conditions they incorporate. Semantic conditions, including organ and tumor maps, can be effectively harnessed to generate images that adhere to specific locality constraints [66, 65, 29, 24, 68]. Textual conditions like medical prompts, which are semantically rich and descriptive, can also be employed to enhance the diversity of generated images [7, 4, 9, 71, 30, 40]. However, despite their advancements, these methods fail to leverage the merits of both types of constraints, resulting in the semantic-guided models being not versatile enough to synthesize samples to a greater diversity and text-guided models not accurate enough to fully reflect exact spatial relationships among anatomical structures. Realizing this disadvantage, MedSyn [64] proposes to adopt an text-guided anatomy-aware generation pipeline that generates CT images not only conforms to anatomical masks, but also to a free-text textual guidance on a specific pulmonary disease. However, MedSyn requires paired semantics and texts as conditional input, and it only synthesizes a local region of the body (chest), which greatly limits its versatility and downstream usability. Moreover, contrary to the field of 2D medical image analysis, how to effectively incorporate textual knowledge and align it with semantic information remains largely underexplored in 3D CT synthesis [23]. This deficiency of targeted researches significantly hinders the formulation of an end-to-end, i.e. text to samples of masks and images, generation framework of medical CT dataset.
To overcome these limitations, we propose a novel framework GuideGen that synthesizes full-torso anatomies and CT images based on a single medical free-text input, which are constructed by incorporating patient demographic information and clinical knowledge for a potential tumor site. Naturally, GuideGen can synthesize paired samples that provides for dataset construction on downstream multi-organ and tumor segmentation tasks. It is also more user-friendly as it necessitates only an medical free-text input, empowering researchers to synthesize full-torso CT dataset with minor effort. Our GuideGen consist of three individual stages, separately for text-guided semantics synthesis, contrast-aware CT autoencoding and latent-guided CT generation. In Stage-I, we offer a text-conditional semantic synthesizer based on a categorical diffusion model to generate exact discrete label indices, unlike current practices [24] that suffer from an ambiguity problem (See Sec. 3.1). It also translates the implicit information in medical prompts onto concrete semantic conditions to better guide the subsequent CT generation. In Stage-II, we train a contrast-aware autoencoder to extract comprehensive anatomical features from multiple contrast levels, avoiding the detail degradation by truncating intensities to a specific range that favors a region of interest as per many current practices [8, 68]. In Stage-III, we utilize a diffusion latent generator that operates on the combined information of semantic latents derived from anatomies in Stage-I, image latents extracted in Stage-II and textual latents capturing demographic or clinical features in the input medical prompt.
We train and validate GuideGen on an assembly of 12 public tumor datasets and one in-house colorectal cancer dataset, which contain the multi-modal information of CT images and medical descriptions. To testify GuideGen’s downstream usability, we further include 2 multi-organ segmentation datasets and 3 tumor segmentation datasets. Our framework achieves state-of-the-art results in sample quality, conditional consistency as well as downstream usability on segmentation tasks across multiple datasets.
Our contributions can be concluded as follows:
  • We propose GuideGen, a viable and novel solution capable of performing text-based full-torso anatomy and CT generation.
  • We introduce a text-conditional semantic synthesizer that avoids ambiguity in current 3D anatomy generations, a contrast-aware module to accommodate high intensity variations in full-torso CT volumes and a latent-guidance module to better preserve detailed semantic guidance.
  • We conduct thorough experiments comparing the samples generated by GuideGen with 6 state-of-the-art generative frameworks from their quality, conditional consistency and downstream usability, on all of which GuideGen achieves better performances.
2 Related Works
Generative frameworks
such as Energy-Based Models [17, 21], Generative Adversarial Networks [15, 61, 73], Normalizing Flows [22, 36], Variational Autoencoders [41], and Diffusion Models [67, 32, 33] encouraged myriad researches on image generation due to their high authenticity and variability. Among these studies, the most advantageous feature that has emerged is arguably the capacity to steer the generation process via user-defined conditions [56]. This capability allows for a high degree of precision and customization in creating images, responding directly to the specific requirements set by the user, while providing a way for the generated images to be used for dataset augmentation purposes [19, 35]. In this work, we extends the ability of current diffusion models for a text-guided anatomy and CT synthesis, which mitigates data scarcity often experienced in the field of medical image analysis.
Medical CT Synthesis
on which previous endeavors undoubtedly provide some performance boost on their tasks of choice [25, 39, 13, 45, 29]. Despite their pioneering efforts, their work is confined to specific organs, which renders repeatedly retraining the network a necessity for each specific organ of interest. A lack of external guidance also underscores the need for a more flexible approach that conforms to not only descriptive texts but also to detailed semantics. Although recent work has begun to focus on these problems [64, 40], their generation processes require paired semantic and textual input to generate regional patches, which limits its downstream applicability. While a step forward, this indicates room for further development to achieve a more automated and generalized solution. This paper intends to provide a text-guided full-torso generative framework that aids to a wide range of downstream tasks.
3 Method
As illustrated in Fig. 1, GuideGen differs from current text-conditional medical image synthesizers [7, 23, 64] in that it disentangles the generative process into three stages separately for text-based anatomy placement, contrast-aware feature extraction and latent-guided image generation to address the multitude of labels, diversity of intensities and the incorporation of textual constraints in full-torso anatomy and CT volumes. We will elaborate on the specifics of our designs in the following sections.
3.1 Stage-I: Text-conditional Semantic Synthesizer
Current generative frameworks that uses dedicated semantic synthesizers [24, 10] unintentionally introduce ambiguities as they struggle to capture the sharp transitions in labels near semantic boundaries due to an inaccurate data modeling. They also fail to provide an interface for more fine-grained and versatile control via free-form texts. Contrary to these methods, our Text-Conditional Semantics Synthesizer (TCSS) provides a text-based ambiguity-free solution by building upon a categorical diffusion model [28, 70] which directly assumes a discrete formulation. As shown in Fig. 1(a), at its core, we train a categorical diffusion $p_{\theta}$ parameterized by $\theta$ that models the underlying distribution of mask volumes $\mathbf{m}$. At the first timestep, the diffusion variable $\mathbf{x}_{0}=\mathbf{m}\in\{1,\cdots,N\}^{H\times W\times D}$, where $H$, $W$, $D$ separately denotes mask height, width and depth. In the forward process, $\mathbf{x}_{0}$ is gradually transformed to a categorical noise $\mathbf{x}_{T}\sim\mathcal{C}_{N}(\mathbf{x}_{T};\mathbf{1}/N)$ in $T$ timesteps under a noise schedule $\beta_{1:T}$, where $\mathcal{C}_{N}$ denotes a categorical distribution of $N$ categories (semantic classes). The forward probability transition can be described by
$q(\mathbf{x}_{t}^{i}|\mathbf{x}^{i}_{t-1})=\mathcal{C}_{N}(\mathbf{x}_{t}^{i};(1-\beta_{t})\mathbf{e}(\mathbf{x}^{i}_{t-1})+\beta_{t}\cdot\frac{\mathbf{1}}{N})\in[0,1]^{N},$ (1)
where the superscript $i$ is the voxel index and $\mathbf{e}(\cdot)$ is a function that returns one-hot probability vectors from a categorical input. For compactness and ease of reading, from now on we will consider the diffusion process on an image with only one voxel to lose the superscript $i$, i.e. $H=W=D=1$. Supposing voxels are independent, the formula derived below can be safely extended to images of any size. From the properties of Markov chains [28], the forward process of $p_{\theta}$ can be described as
$q(\mathbf{x}_{t}|\mathbf{x}_{0})=\mathcal{C}_{N}(\mathbf{x}_{t};\bar{\alpha}_{t}\mathbf{e}(\mathbf{x}_{0})+(1-\bar{\alpha}_{t})\cdot\frac{\mathbf{1}}{N})\in[0,1]^{N},$ (2)
where $\bar{\alpha}_{t}=\prod_{\tau=1}^{t}(1-\beta_{\tau})$.
To extract relevant information in the medical prompt for semantic placement, we use a text encoder $\mathcal{E}_{T}$ pretrained on medical knowledge [62] to map the original prompt onto a latent space and perform cross-attention to guide anatomical generation by extracting spatial relation among anatomies. We interleave these cross-attention layers with a 3D U-Net [11] backbone to construct $p_{\theta}$. Specifically, for each intermediate layer $l$ in the 3D U-Net, the latent vector $\mathbf{z}_{l}$ is first computed by
$\mathcal{Q}=W^{q}(\mathbf{z}_{l-1}),\mathcal{K}=W^{k}(\mathcal{E}_{T}(\mathbf{p})),\mathcal{V}=W^{v}(\mathcal{E}_{T}(\mathbf{p}))$ (3)
$\mathbf{z}_{l}=\text{MLP}(\text{LayerNorm}(\text{Softmax}(\mathcal{Q}\mathcal{K}^{\top}/\sqrt{d})\mathcal{V})),$ (4)
where $W^{q}$, $W^{k}$, and $W^{v}$ are learnable projection mappings, $\mathbf{p}$ is the input medical prompt and $d$ is the context dimension. We train $p_{\theta}:[0,1]^{N}\rightarrow[0,1]^{N}$ by minimizing the Kullback-Leiber divergence between posterior and generator distributions across the diffusion process as in [27]. By denoting the posterior $q(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0})=\frac{q(\mathbf{x}_{t}|\mathbf{x}_{t-1})q(\mathbf{x}_{t-1}|\mathbf{x}_{0})}{q(\mathbf{x}_{t}|\mathbf{x}_{0})}\in[0,1]^{N}$ as $\boldsymbol{\pi}_{t}(\mathbf{x}_{0})$, the overall training objective is
$\mathcal{L}=\mathbb{E}_{t,\mathbf{x}_{0}}\left[\mathcal{D}_{KL}\left(\mathcal{C}_{N}(\mathbf{x}_{t-1};\boldsymbol{\pi}_{t}(\mathbf{x}_{0}))||\mathcal{C}_{N}(\hat{\mathbf{x}}_{t-1};p_{\theta})\right)\right],$ (5)
with the hat notation distinguishing generated variables.
To avoid training instability, we make a similar reparameterization as in DDPM [27] for enhanced performance by training a generator model on the same space as the input $\mathbf{x}_{0}$. Instead of training a model $p_{\theta}(\mathbf{x}_{t},\mathbf{p})$ that models on the distribution $\mathcal{C}_{N}(\hat{\mathbf{x}}_{t-1};p_{\theta})$, we will train and draw samples from its reparameterized version, denoted as $f_{\theta}=\mathcal{C}_{N}(\hat{\mathbf{x}}_{0};f_{\theta}(\mathbf{x}_{t},\mathbf{p}))$. The training loss $\mathcal{L}$ needs to be modified accordingly: From the total probability formula, denoting $\mathbf{A}=(\boldsymbol{\pi}_{t}(1),\cdots,\boldsymbol{\pi}_{t}(N))$, we have $p_{\theta}=\mathbf{A}f_{\theta}$. The voxel-wise loss for TCSS can be written as
=$\begin{split}\mathcal{L}&{}_{TCSS}=\\ &\mathbb{E}_{t,\mathbf{x}_{0}}\left[\mathcal{D}_{KL}\left(\mathcal{C}_{N}(\mathbf{x}_{t-1};\boldsymbol{\pi}_{t}(\mathbf{x}_{0}))||\mathcal{C}_{N}(\hat{\mathbf{x}}_{t-1};\mathbf{A}f_{\theta})\right)\right].\end{split}$ (6)
Under real settings where an image actually contains multiple voxels, the total loss $\mathcal{L}_{TCSS}$ is averaged over each individual voxel.
At inference, TCSS samples from the distribution $f_{\theta}$ and computes the next denoising step by sampling on the generator distribution $S(\mathbf{A}f_{\theta}(\hat{\mathbf{x}}_{t},\mathbf{p}))$, where $S$ is the sampling scheme. Similar to [70], we sample by probability for timesteps $2\sim T$ and take the index with max probability on the final denoising step to add stability to the generation process.
3.2 Stage-II: Contrast-aware Autoencoder
Similar to [9, 23], we use a pair of image autoencoder to extract high-level texture details from CT volumes while lowering memory demands to incorporate larger CT volumes. As shown in Fig. 1(b), similar to physicians examining different anatomical structures using different intensity windows, our Contrast-aware Autoencoder (CAE) strives to accurately accommodate the entire dynamic range of CT intensities with the help of a contrast-aware rescaler $h$. Formally,
$h(\mathbf{x})=k\max\left\{\min\left\{\frac{\mathbf{x}-w_{c}+w_{r}}{2w_{r}},1\right\},0\right\}+b,$ (7)
where $w_{c}$ and $w_{r}$ are separately known as the window center and window radius. By definition, $h$ truncates the intensity range of $\mathbf{x}$ to $[w_{c}-w_{r},w_{c}+w_{r}]$ and rescales them to $[b,k+b]$. At each training step, we randomly sample the window center $w_{c}$ and window radius $w_{r}$ from the intensity range of $\mathbf{x}$ and use learnable $k$ and $b$ to map the truncated result back to the input space. Separately denoting the encoder and decoder halves of CAE as $\mathcal{E}_{I}$ and $\mathcal{D}_{I}$, the $l_{1}$ reconstruction loss on an input volume $\mathbf{v}$ can be defined as
$\mathcal{L}_{rec}=||h(\mathcal{D}_{I}(\mathcal{E}_{I}(\mathbf{v})))-h(\mathbf{v})||.$ (8)
Aside from $\mathcal{L}_{rec}$, we follow the scheme of VQGAN [18] to introduce a quantization step $\mathbf{q}$ with its adversarial components, learnable codebook $\mathbf{c}$ and corresponding losses $\mathcal{L}_{vq}$ to the autoencoder training process. Specifically, CAE encodes an input volume $\mathbf{v}$ to its latent representation $\mathcal{E}_{I}(\mathbf{v})$, replaces each feature $\mathbf{z}$ with its nearest codebook entry $\mathbf{q}(\mathbf{z})=\mathbf{c}_{\mathbf{z}}$, and tries to recover the original volume $\mathbf{v}$ from the quantized latent vectors. For adversarial losses, we use a frame discriminator $\mathbf{D}_{f}$ that randomly chooses several planar slices from the CT volume and tries to distinguish between authentic slices and synthetic ones. Aside from this, we introduce a volume discriminator, $\mathbf{D}_{v}$, that does the same job from a volumetric viewpoint. These discriminators, separately focusing on anatomical details within a CT slice and the general structure of a CT volume, are not subject to contrast rescaling to avoid training instability and preserve the range of intensities for each organ. We also use perceptual loss based on a VGG-16 network [72, 59] $\mathcal{L}_{perc}$ for reducing artifacts in image reconstruction as in [8]. Overall, the loss function for CAE is
$\begin{split}\mathcal{L}_{CAE}=&\mathcal{L}_{rec}(h(\mathcal{D}_{I}(\mathcal{E}_{I}(\mathbf{v}))),h(\mathbf{v}))+\\ &\mathcal{L}_{vq}(\mathcal{E}_{I}(\mathbf{v}),\mathbf{c})+\\ &\mathcal{L}_{perc}(h(\mathcal{D}_{I}(\mathcal{E}_{I}(\mathbf{v}))),h(\mathbf{v}))+\\ &\mathcal{L}_{disc}^{frame}(\mathcal{D}_{I}(\mathcal{E}_{I}(\mathbf{v}))[i],\mathbf{v}[i],\mathbf{D}_{f})+\\ &\mathcal{L}_{disc}^{volume}(\mathcal{D}_{I}(\mathcal{E}_{I}(\mathbf{v})),\mathbf{v},\mathbf{D}_{v}),\end{split}$ (9)
where $\mathcal{L}_{vq}(\mathcal{E}_{I}(\mathbf{x}),\mathbf{c})=||\Omega(\mathcal{E}_{I}(\mathbf{x}))-\mathbf{c}_{\mathbf{z}}||_{2}^{2}+\alpha||\Omega(\mathbf{c}_{\mathbf{z}})-\mathcal{E}_{I}(\mathbf{x})||_{2}^{2}$ and $\mathcal{L}_{disc}^{\cdot}(\mathbf{x}^{\prime},\mathbf{x},\mathbf{D})=\log\mathbf{D}(\mathbf{x})+\log(1-\mathbf{D}(\mathbf{x}^{\prime}))$. $\Omega$ here is the stop-gradient operation.
3.3 Stage-III: Latent-guided Feature Generator
Since the encoded image and its semantic condition are not of the same shape, current practices usually downsample the mask volume for them to be consistent [8]. However, spatial relations between tumor and organs can be lost if the downsampling factor is too high. To promote preservation of semantic details and the textual guidance from medical prompts, we propose a Latent-guided Feature Generator (LFG) that operates entirely on the latent space to generate fine-grained image latents. As shown in Fig. 1(c), firstly, LFG encodes the torso semantics $\mathbf{m}_{A}$ and tumor semantics $\mathbf{m}_{T}$ divided from anatomies generated in Stage-I by several strided convolutional layers to downsample their spatial size. Between each downsampling convolution step, their semantic latents $\mathbf{z}^{A}$ and $\mathbf{z}^{T}$ are fused by attention layers
$\begin{split}\mathbf{z}^{A}&=\text{MLP}(\text{LayerNorm}(\text{Softmax}(\mathcal{Q}\mathcal{K}^{\top}/\sqrt{d})\mathcal{V}))\\ \end{split}$ (10)
where $\mathcal{Q}=W^{q}(\mathbf{z}^{A})$, $\mathcal{K}=W^{k}(\mathbf{z}^{T})$, $\mathcal{V}=W^{v}(\mathbf{z}^{T})$ are attention matrices and $W^{q}$, $W^{k}$, $W^{v}$ are learnable mappings. Secondly, we use a diffusion model to approximate the distribution of image latents $\mathbf{z}_{0}$. In the forward process, $\mathbf{z}_{0}$ is gradually converted to Gaussian noise $\mathbf{z}_{T}\sim\mathcal{N}(0,\mathbf{I})$ through $T$ timesteps following a noise schedule $\beta^{\prime}_{1:T}$
$q(\mathbf{z}_{t}|\mathbf{z}_{t-1})=\mathcal{N}(\mathbf{z}_{t};\sqrt{1-\beta^{\prime}_{t}}\mathbf{z}_{t-1},\beta^{\prime}_{t}\mathbf{I}).$ (11)
We follow a similar way as in Eq. 4 to use cross-attention layers to inject textual latents into our diffusion backbone and concatenate semantic latents $\mathbf{z}^{A}$, $\mathbf{z}^{T}$ to the diffusion variable $\mathbf{z}_{t}$ at each timestep. Finally, we construct the generative objective as in [27]:
$\mathcal{L}_{LFG}=\mathbb{E}_{t,\boldsymbol{\epsilon}\sim\mathcal{N}(0,\mathbf{I})}\left[||\boldsymbol{\epsilon}-f_{\varphi}(\mathbf{z}_{t},\mathbf{z}^{A},\mathbf{z}^{T},\mathcal{E}_{T}(\mathbf{p}))||_{2}^{2}\right],$ (12)
where $f_{\varphi}$ is the 3D U-Net [11] diffusion backbone parameterized by $\varphi$ and $\mathcal{E}_{T}(\mathbf{p})$ is the textual latents encoded by the same text encoder described in Sec. 3.1. At inference time, $f_{\varphi}$ gradually recovers an image latent $\hat{\mathbf{z}}_{0}$ by reverse sampling from random noise, which is then decoded into a valid CT image by the decoder $\mathcal{D}_{I}$ from CAE.
4 Experiments
Dataset Construction
(1) Training: We compile multiple CT datasets to train GuideGen, including 12 public imaging datasets with their corresponding clinical information from The Cancer Imaging Archive (TCIA) [12] and one real-world, private dataset (RJ) on colorectal cancer. Their groundtruth label maps are retrieved using pretrained multi-organ and tumor segmentation networks [63, 34]. This collected dataset is divided into 4534 training and 1179 validation cases used for generative training. (2) Inference: For downstream usability evaluation, we additionally incorporate BTCV [46] and AMOS22 [37] datasets for multi-organ segmentation tasks and MSD-Lung Tumors (LU), MSD-Colon Cancer (CO) [3], KiTS21 [26] (KI) datasets for tumor segmentation. The rest of the experiments are conducted on the validation split of TCIA and RJ. We will elaborate on our experimental settings in Appendix 6.1.
Evaluation Metrics
The quality of our generated images are evaluated through the Learned Perceptual Image Patch Similarity (LPIPS), Fréchet Inception Distance (FID) and Fréchet Video Inception Distance (FVD) as per the general paradigm [23, 69]. We also use some these metrics to evaluate the quality of our masks due to a lack of dedicated metrics. Additionally, we report the Dice Similarity Coefficient (DSC) to evaluate image-mask consistency and the accuracy of several dedicated classifiers to measure alignment between image and certain feature described in the input prompt. Downstream Usability is evaluated through DSC and 95% Hausdorff Distance (HD95) as commonly used in segmentation tasks [14, 38]. We set HD95 to 1000 to avoid NaN values for valid comparison when there are no semantic predictions. We calculate the standard deviation for usability metrics and use a double sided $t$-test to determine the statistical significance of GuideGen’s improvement over its nearest competitor and report them in Appendix 6.2.
Implementation Details
All generative trainings are performed under an initial learning rate of $1\times 10^{-4}$ with a batch size of 1 per GPU on a multi-GPU(4) server and a patch size of $128\times 128\times 128$. The training processes use AdamW [49] optimizers with a momentum of 0.99, a weight decay of $1\times 10^{-5}$ and a linear learning rate scheduler to gradually decrease the learning rate to 0 over 500 epochs. For each diffusion model in TCSS and LFG, we use cosine noise schedule with 1000 timesteps and LFG uses DDIM denoising [60] of 200 timesteps to recover CT latents from Gaussian noise at a guidance scale of 2.5. Moreover, our CAE downsamples the original image by 4×4$\times$4 × and uses a codebook of 8192 entries.. At inference time, we train several segmentation networks using the official implementation of nnU-Net [34]. GuideGen itself is implemented using PyTorch [53] and all experiments are performed on 4 NVIDIA 4090 GPUs.
4.1 Main Results
Generation quality
Methods infer. model comput.cost Full Anatomy Tumor Only
time(s) param.(M) (GFLOPs) LPIPS FID LPIPS FID
MedGen3D [24] 67.2 48.8 1295.7 0.70 201 0.29 33.5
LDM [57] 19.9 115.2 314.9 0.67 98.6 0.30 69.1
Hu’s [29] 0.8 - 0.15 - - 0.32 64.8
GuideGen 54.1 49.1 2842.3 0.39 11.0 0.35 32.1
Table 1: Comparison studies of different semantic generation methods of full-torso anatomies and tumor-only binary masks.
        Methods         input condition         inference         model         computational         generative metrics
        mask         text         time(s)         parameters(M)         cost(GFLOPs)         LPIPS         FID         FVD128
        Pinaya’s [54]         N         Y         70.2         196.1         441.0         0.465         92.0         2091
        GenerateCT[23]         N         Y         102.0         288.7         2167.9         0.521         151.7         4634
        MONAI-GM [55]         G         N         74.1         166.5         461.9         0.406         57.7         1890
        Zhuang’s [74]         G         N         45.5         81.2         2752.2         0.327         23.2         1094
        MedDDPM [16]         G         N         42.4         80.5         2748.8         0.354         35.6         1299
        MedSyn [64]         G         Y         24.7         115.1         1832.6         0.463         27.9         2267
        GuideGen         G         Y         26.1         196.8         491.4         0.242         20.5         785
        MONAI-GM [55]         R         N         74.1         166.5         461.9         0.393         54.6         1791
        Zhuang’s [74]         R         N         45.5         81.2         2752.2         0.335         39.1         1366
        MedDDPM [16]         R         N         42.4         80.5         2748.8         0.342         45.8         1619
        MedSyn [64]         R         Y         24.7         115.1         1832.6         0.396         50.1         1960
        GuideGen         R         Y         26.1         196.8         491.4         0.237         20.7         714
Table 2: Comparison studies of different image generation frameworks. The “$128$” superscript on FVD metric indicates we specifically consider CT volumes of 128 slices. A mask condition of ‘G’ or ‘R’ separately denotes semantic guidance from GuideGen-generated masks and real masks.
(1) Mask quality: To validate that GuideGen’s discrete formulation is indeed superior to current methods based on continuous modeling, we choose MedGen3D [24], which is based on DDPM [27], and LatentDiffusion [57] as baseline methods and infuse them with textual information in the same way as ours. Their generation results are rounded to the nearest integer to represent semantic information. We also compare with the Hu’s deformation method described in [29] for tumor generation. The results are shown in Tab. 1. It is evident that GuideGen performs better to baselines by a large margin on anatomy generation. This is not surprising as the ambiguity problem mentioned in Sec. 3.1 significantly hinders the performance of MedGen3D and LDM when the number of semantic classes is large. This problem persists but is less prominent as the number of semantic classes drops, as shown by synthesizing only the tumor masks, where GuideGen maintains comparable performance to comparison methods. (2) Image quality: We compare GuideGen’s image generation quality with a series of generative methods from different organs, including Pinaya’s and MedDDPM [54, 16] for brain MRI generation, GenerateCT [23] and MedSyn [64] for lung CT generation, Zhuang’s [74] for abdominal CT generation and the generic medical generative framework MONAI-GM [55]. All of them are trained using the train split from TCIA and RJ datasets. At inference time, for methods that dictates a mask input, we compare their performances with GuideGen’s performance under both GuideGen-generated and real masks to ensure a fair comparison, as shown in Tab. 2. We observe that GuideGen consistently outperforms existing methods in image quality under all metrics. Particularly, we notice that methods based solely on text conditions usually perform worse than those on semantic masks. We speculate this is because the highly diverse CT textures and intensity ranges among organs are difficult to recover without voxel-wise guidance, which reflects the necessity of our TCSS. Moreover, we see little discrepancy between metrics computed from samples generated using GuideGen-generated masks and real-world masks, reflecting that our TCSS can provide comparable quality to real masks. We also evaluate the generation results qualitatively as shown in Fig. 2 by observing that existing methods do not adapt well to the high dynamic range of CT intensities, appearing either too bright or too dark in specific body regions. In contrast, GuideGen maintains a realistic CT intensity range across the full torso, thanks to our proposed CAE.
Generation-condition alignment
Methods DSC
Spl. Kid. Liver Sto. Pan. Lung S.B. Duo. Colon Heart Avg.
MONAI-GM 0.73 0.72 0.80 0.60 0.43 0.84 0.49 0.35 0.55 0.40 0.59
Zhuang’s 0.43 0.43 0.62 0.36 0.21 0.69 0.37 0.23 0.26 0.24 0.38
MedDDPM 0.35 0.36 0.39 0.54 0.29 0.34 0.47 0.43 0.67 0.22 0.41
MedSyn 0.52 0.51 0.51 0.40 0.07 0.59 0.12 0.01 0.54 0.22 0.35
GuideGen 0.73 0.73 0.81 0.60 0.42 0.83 0.57 0.37 0.60 0.41 0.61
Table 3: Image-mask alignment between generated CT and their full-torso anatomical guidance. We report the DSC scores on ten major thoracic and abdominal organs.
Methods Accuracy
Age Gender Race Tumor Loc. Avg.
Pinaya’s 0.06 0.35 0.10 0.17 0.17
GenerateCT 0.07 0.21 0.44 0.03 0.19
MedSyn 0.17 0.74 0.51 0.47 0.47
GuideGen 0.46 0.88 0.57 0.94 0.71
Table 4: Image-text alignment between generated CT and textual guidance. We report the accuracy of four dedicated classifiers on demographics and clinical features derived from the given medical prompts.
Figure 2: Qualitative results of different generation methods conditioned on the same textual prompts. Mask inputs to baseline models are generated with GuideGen. Two slices are shown per case. For better visualization, we use a CT intensity window of [-975,-225]HU for displaying the chest region (rows 1-2), and [-50,150]HU for the abdominal region (rows 3-6) [58]. “T” and “S” stands for frameworks that requires textual inputs or semantic inputs, respectively. More qualititative results can be found in Appendix 6.2.
      Dataset       Method       DSC
      Spleen       Kidneys       Liver       Sto.       Pan.       Adr.       Eso.       Aorta       IVC       Gall.       Duo.       Blad.       PV&SV       Avg.
      BTCV       Real       0.92       0.79       0.94       0.86       0.7       0.6       0.71       0.89       0.81       0.52       -       -       0.52       0.74
      5pt.       MONAI-GM       0.91       0.89       0.94       0.80       0.61       0.44       0.60       0.84       0.78       0.29       -       -       0.48       0.69
      Zhuang’s       0.90       0.88       0.95       0.83       0.65       0.54       0.69       0.88       0.82       0.49       -       -       0.56       0.74
      MedDDPM       0.92       0.90       0.95       0.87       0.66       0.54       0.68       0.88       0.84       0.39       -       -       0.49       0.74
      MedSyn       0.89       0.90       0.96       0.81       0.65       0.56       0.70       0.86       0.86       0.39       -       -       0.32       0.72
      GuideGen       0.94       0.92       0.96       0.91       0.74       0.59       0.73       0.91       0.87       0.49       -       -       0.65       0.79
      AMOS       Real       0.95       0.94       0.96       0.89       0.81       0.67       0.78       0.92       0.87       0.72       0.76       0.82       -       0.84
      5pt.       MONAI-GM       0.83       0.84       0.91       0.74       0.60       0.50       0.60       0.81       0.71       0.34       0.53       0.43       -       0.65
      Zhuang’s       0.85       0.87       0.90       0.66       0.63       0.46       0.61       0.82       0.73       0.26       0.55       0.62       -       0.66
      MedDDPM       0.86       0.89       0.90       0.74       0.61       0.46       0.61       0.86       0.76       0.43       0.55       0.60       -       0.69
      MedSyn       0.85       0.90       0.91       0.70       0.64       0.52       0.62       0.80       0.69       0.21       0.53       0.64       -       0.67
      GuideGen       0.93       0.92       0.95       0.86       0.75       0.55       0.73       0.88       0.81       0.65       0.67       0.73       -       0.79
Table 5: Segmentation performance on multi-organ segmentation tasks using real or synthetic multi-organ masks from each generative framework.
We will thoroughly evaluate GuideGen’s conditional alignment from three aspects while comparing with other baselines. (1) Image-mask alignment: Image-mask alignment can be quantitatively measured by comparing predicted results from pretrained segmentation models [63] on generated images with the actual masks given, with a higher alignment being a higher segmentation accuracy. We report the DSC of 10 major organs between predicted masks and real semantic condition input in Tab. 3. It can be seen that GuideGen behaves desirably for semantic correspondence, achieving a 2% gain on DSC compared to its nearest competitor. (2) Image-text alignment: To quantitatively evaluate the alignment between generated images and text prompt conditions, similar to [20, 23], we use a series of pretrained classifiers on certain features (age, gender, race, tumor location) to judge whether the generated images faithfully exhibits features specified in medical prompts. As shown in Tab. 4, images generated by our framework faithfully reflect the gender information and tumor location while outperforming other text-conditioned frameworks in preserving age-related and racial features. (3) Mask-prompt alignment: We measure mask-prompt alignment only qualitatively since little information except spatial relationships is depicted in generated masks. We give visualizations of mask-prompt alignment using mask synthesizers in MedGen3D (adapted by adding textual constraints to generation backbone as in Eq. 4) and GuideGen in Fig. 3 by varying tumor locations in our text prompt. It can be seen that our discrete formulation not only generates masks of accurate shape and structure, but also better preserves the spatial relationships between tumor and organs.
Usability in segmentation dataset synthesis
To assess the usability in synthesizing datasets for segmentation purposes, we use nnU-Net [34] to train segmentation models using samples generated by different frameworks as well as real-world samples from established downstream datasets. Sample synthesis is based on the anatomical masks and/or textual prompts in the validation split of our TCIA and RJ datasets.
Figure 3: Qualitative results of generated full-torso anatomical masks, with tumor masks masked in red. For each textual constraint specified, GuideGen can provide faithful anatomies while being visibly superior thanks to our discrete formulation.
(1) Multi-organ segmentation: As shown in Tab. 4.1, segmentation models trained on samples generated from GuideGen can achieve comparable performance as those trained on real samples while outperforming those trained using synthetic samples generated with baseline frameworks, highlighting the generation quality of our generalized framework. (2) Tumor segmentation: Results for tumor segmentation (Tab. 6) show that our model can generate samples of lung tumors competitive with real ones while being more beneficial for other tumor segmentation tasks than baseline methods. Qualitative results can be found in Appendix 6.2.
4.2 Ablation studies
      Method       DSC       $\text{HD}_{95}$
      LU       CO       KI       Avg.       LU       CO       KI       Avg.
      Real       0.69       0.47       0.72       0.63       11.9       212.7       70.9       98.5
      MONAI-GM       0.48       0.10       0.24       0.27       31.7       284.9       293.5       203.4
      Zhuang’s       0.12       0.07       0.13       0.11       716.0       274.1       511.5       500.5
      MedDDPM       0.10       0.09       0.29       0.16       776.0       273.6       119.4       389.7
      MedSyn       0.44       0.11       0.39       0.31       33.1       267.4       309.0       203.2
      GuideGen       0.69       0.24       0.51       0.48       10.5       244.0       154.5       136.3
Table 6: Segmentation performance on tumor segmentation tasks of different organs using real or synthetic cases from each method.
         Mask Synthesizers          LPIPS          FID          FVD128
         None          0.56          279          5981
         LDM          0.40          61.4          2217
         MedGen3D          0.41          69.6          2045
         TCSS          0.24          20.5          785
Table 7: CT generation quality with different mask synthesizers. “None” denotes that the mask condition is removed (only the text condition is used).
GuideGen Settings generative metrics usability metrics
CAE LFG LPIPS FID FVD128 Avg. DSC Avg. HD95
N N 0.35 48.1 1729 0.29 201.9
N Y 0.36 46.4 1476 0.30 215.1
Y N 0.25 23.1 1004 0.37 180.4
Y Y 0.24 20.5 785 0.48 136.3
Table 8: Ablation studies on CAE and LFG modules. We use “Y”/“N” in CAE to indicate whether a contrast-aware rescaler is in effect and “Y”/“N” in LFG to indicate whether we use cross-attention in semantic guidance described in Sec. 3.3 as opposed to direct mask downsampling as in [8].
We evaluate the effectiveness of each component in GuideGen, including the semantic synthesizer in TCSS, contrast-aware module in CAE and the efficacy of LFG as opposed to using latent diffusion models with straightforward mask downsampling. The usability metrics are computed in the same way as Tab. 6. From Tab. 7, we observe that CT synthesized using TCSS-generated semantics performs the best, and a lack of semantic guidance degrades performance drastically. Comparing Tab. 7 with Tab. 1, we can see an interesting trend that generated image quality is roughly in line with the quality of generated masks. Tab. 8 shows a clear boost in generative metrics by integrating contrast-awareness to image autoencoders for full-torso CT generation. The LFG strategy also provides for the image quality, and more importantly it benefits in preserving tumor semantics, as shown by the increase in usability metrics. Similarly, CAE also boosts performance in the downstream usability, we suspect this comes as a collateral benefit brought by an increased quality of CT synthesis.
5 Conclusion
In this paper, we have demonstrated a novel framework for generating paired 3D anatomies and 3D CT images covering the entire torso, conditioned on a free-text medical prompt. Firstly, we introduce a text-conditional semantic synthesizer with categorical modeling to aid in the text-based synthesis of semantic masks. Secondly, we propose a contrast-aware autoencoder to accommodate the high dynamic range of CT images. Finally, we propose a latent-guided feature generator to promote detail preservation and textual alignment over conventional resampling methods. Our model outperforms a range of existing methods under comprehensive evaluation, including generative quality, alignment with textual and semantic guidance, and usability in training semantic segmentation models. Overall, our approach not only advances the state of the art in text-guided 3D CT generation but also lays a foundation for more robust training dataset synthesis for medical image analysis, making it highly impactful for clinical applications.
Supplementary Material
6 Appendix
Code Disclosure Statement
All source code, include model training, data processing, downstream evaluation and metric computation are contained in the code.zip submitted along with this appendix, and they will be publicly released should our paper be accepted.
6.1 Datasets and Implementation Details
6.1.1 Dataset Details
Training datasets
We compiled 12 TCIA [12] datasets on three common types of cancer (lung, kidney and colon). To supplement the scarcity of publicly available colon cancer multimodal (paired CT and clinical information) dataset, we also include one private real-world dataset, dubbed RJ, for colorectal cancer. The preprocessing of all the collected unstandardized samples include the following four steps: (a) Data cleansing: All of the collected CT samples are reoriented into RAS orientation and resampled to isotopic spacing, with duplicate, impaired, partial DICOM series identified and removed. (b) Label Extraction: CT images in the collected dataset are segmented by TotalSegmentator [63] and pretrained models [34] on their respective tumor types to retrieve both multi-organ and tumor segmentation masks; their corresponding clinical information are retrieved using a private medical GPT in a certain format: The patient is {demographics group}. In this imaging, the patient’s condition is described as {clinical information}. (c) Noise Removal: Organ labels belonging to the same anatomical class (e.g. individual rib label and the anatomical class ribs) are merged into one semantic label to mitigate noise in the multi-organ labels. Samples with no or erroneous (tumor predictions outside organ prediction) tumor mask predictions are discarded. (d) Dataset Preparation: We randomly split the processed images into 4534 training cases and 1179 validation cases. Specifically, we use the CT, clinical information and masks in the training split to train GuideGen as well as baseline models, detailed description of each training dataset of choice are listed in Tab. 9, and we showcase the distribution of each demographics group in TCIA and RJ datasets in Fig. 4.
       Dataset Name        Description        No. Patients        No. Cases        No. Cases after pp.
       CMB-CRC [5]        Colorectal Cancer        54        237        50
       TCGA-READ [43]        Rectum Cancer        3        34        4
       TCGA-COAD [44]        Colon Cancer        25        93        21
       RJ        Colorectal Cancer        4005        4418        3235
       CPTAC-CCRCC [50]        Renal Cancer        262        727        248
       TCGA-KICH [48]        Renal Cancer        15        109        26
       TCGA-KIRC [1]        Renal Cancer        267        2654        734
       TCGA-KIRP [47]        Renal Cancer        33        376        56
       CPTAC-LSCC [52]        Pulmonary Cancer        212        238        236
       CPTAC-LUAD [51]        Pulmonary Cancer        244        242        71
       CMB-LCA [6]        Pulmonary Cancer        97        696        64
       TCGA-LUAD [2]        Pulmonary Cancer        69        624        120
       TCGA-LUSC [42]        Pulmonary Cancer        37        279        92
Table 9: Datasets used for generative training, quality evaluation and alignment measurement. “RJ” is our private dataset on colorectal cancer while “No. Cases after pp.” is short for “The number of cases after preprocessing”.
Figure 4: Demographics distribution for TCIA and RJ datasets used for generative training as well as quality and alignment evaluations.
Inference datasets
We we choose the abdominal split of Multi-Atlas Labeling Beyond the Cranial Vault (BTCV) [46] and Abdominal [37] Multi-Organ Segmentation (AMOS) datasets and evaluate the segmentation performance on some of their organ labels that overlap with our interest. Additionally, for tumor segmentation, we select two Medical Segmentation Decathlon (MSD) [3] datasets along with the Kidney Tumor Segmentation (KiTS) [26] dataset, separately for pulmonary, colorectal and renal cancer. Their details are documented in Tab. 10. We also downsamples the training cases in each inference dataset to $128\times 128\times 128$ to ensure a fair comparison with the $128\times 128\times 128$ volumes generated by GuideGen.
      Dataset Name       Description       No. Train cases       No. Inference cases       Semantic Classes Reported
      BTCV [46]       Abdominal Organs       24       6       Spl., Kid., Liver, Sto., Pan., Adr., Eso., Aorta, IVC, Gall., PV&SV
      AMOS [37]       Abdominal Organs       240       60       Spl., Kid., Liver, Sto., Pan., Adr., Eso., Aorta, IVC, Gall., Duo., Blad.
      MSD-Lung [3]       Pulmonary Cancer       40       10       Lung Tumor
      MSD-Colon [3]       Colon Cancer       80       20       Colon Tumor
      KiTS [26]       Renal Cancer       210       90       Kidney Tumor
Table 10: Datasets used for downstream segmentation evaluation. Abbreviations: Spl.: Spleen, Kid.: Kidneys, Sto.: Stomach, Pan.: Pancreas, Adr.: Adrenal Glands, Eso.: Esophagus, IVC: Inferior Vena Cava, Gall.: Gallbladder, PV&SV: Portal Vein and Splenic Vein, Duo.: Duodenum, Blad.: Urinary Bladder.
6.1.2 Implementation Details
GuideGen
In Stage-I, we use TCSS to generate 19 major pleural, abdominal and pelvic structure***The 19 major anatomical structure of choice are: Lung, Heart, Ribs, Vertebra, Esophagus, Aorta, Inferior Vena Cava, Liver, Spleen, Stomach, Duodenum, Gallbladder, Splenic Vein and Portal Vein, Kidneys, Adrenal Glands, Pancreas, Small Bowel, Colon (with rectum) and Urinary Bladder, covering major anatomies throughout the torso. and 1 tumor semantics. The reason for not choosing more organ labels lies largely in that other anatomies suffer from relatively poor multi-organ segmentation accuracies. It should be noted that our framework should be simulate as many anatomies as possible if the quality of training data is ensured. The model channels in each layer of our backbone U-Net network are 32, 64, 128, 256, respectively, with 8$\times$ downsampling at the U-Net bottleneck. Each layer consists of two ResNet convolution blocks and we add a cross-attention layer to the first and third layer to integrate textual information.
In Stage-II, our CAE downsamples the CT image by 4$\times$ and operates on a codebook of 8192 entries, each with a feature dimension of 8. The frame discriminator $\mathbf{D}_{f}$ in CAE chooses 8 slices for discriminative loss computation in each iteration. We also set a min-max normalization for CT images prior to their input into CAE model: CT images are linearly rescaled in such a way that the CT intensities of the range [-1500, 1500]HU resides in the interval [-1, 1] after rescaling. In this way, we can recover CT intensities from the decoder half of CAE after LFG generation by multiplying the result by the multiplier 1500.
In Stage-III, we first encode tumor and organ semantics separately to a vector with a feature dimension of 4 and spatial dimension 4$\times$ smaller than original input (which computes to $32\times 32\times 32$ in our experiment setting). As we jointly generate tumor and organ semantics in the first stage, we follow the following steps to retrieve label for them individually: Firstly, for 19 organ labels $l_{A}=\{l_{A}^{1},\cdots,l_{A}^{19}\}$, 1 tumor label $l_{T}$ and 1 background label $l_{B}$, the final step of TCSS sampling trajectory will output a probability vector $\hat{\mathbf{x}}_{0}\in[0,1]^{21\times H\times W\times D}$. $H$, $W$ and $D$ are spatial dimensions, which all equals 128 in our setting. Secondly, we perform argmax operation on the channel dimension to retrieve these semantic masks, specifically, the organ semantics $\mathbf{m}_{A}$ and tumor semantics $\mathbf{m}_{T}$ are retrieved by
$\mathbf{m}_{A}=\mathop{argmax}_{l_{A},l_{B}}\hat{\mathbf{x}}_{0};\;\mathbf{m}_{T}=\delta(\mathop{argmax}_{l_{A},l_{T},l_{B}}\hat{\mathbf{x}}_{0},l_{T}).$ (13)
That is, we discard the probabilities of tumor label when retrieving organ semantics, and $\delta(x,y)$ here is the impulse function that takes 1 on the entry $i$ when $x[i]$ equals $y$ and 0 otherwise. The encoded semantic latents, along with the image feature encoded by the encoder half of CAE are then concatenated to a latent vector with a feature dimension of 16, which is fed into a backbone U-Net with the same structure as that described in TCSS. Loss is computed between the U-Net output and the image latent.
Baseline models
We follow official implementation available online for each baseline method, and adjust their model parameters (mainly the channel numbers for each convolutional layer) so that they can generate $128\times 128\times 128$ volumes under our physical setting (NVIDIA RTX 4090 GPUs). We also skip the final super-resolution module in some baselines [64, 23] since it has no relevance to conditional generation and can be added to any generative framework as a postprocessing module.
6.2 Supplementary Results
This section provides tables and figures that serves as a supplement to the main results displayed in the main article.
6.2.1 Qualitative Results
In this subsection, we first showcase GuideGen’s generative performance of both unhealthy and healthy samples using randomly chosen demographic and clinical descriptions from TCIA and RJ datasets in Fig. 5. Secondly, we display multi-organ segmentation results and tumor segmentation results separately in Fig. 6 and Fig. 7. For slices located in the pleural region, we display them using a normalization window at [-975, -225]HU, while for slices in the abdominal and pelvic regions, we display them using a normalization window at [-50, 150]HU as recommended in [58].
6.2.2 Quantitative Results
We display standard variances of each metric used in multi-organ segmentation and tumor segmentation evaluations separately in Tab. 6.2.3 and Tab. 12 in this subsection. We also report the validation accuracy and AUROC of demographic or clinical classifiers used to assess image-prompt alignment in Tab. 13. They are implemented based on the encoder half of U-Net 3D networks [11] and trained on the training split of TCIA and RJ datasets. Classifier classes are chosen following the demographic classes in Fig. 4 and tumor locations in Tab. 9. Metrics in Tab. 13 are computed on the validation split of TCIA and RJ datasets.
6.2.3 Utilizing GuideGen for Dataset Synthesis or Augmentation
Finally, in Tab. 14, we tested the usability of generated samples by using different number of GuideGen-generated cases for dataset synthesis (segmentation models trained with no real cases required) or dataset augmentation (using GuideGen-generated cases as a complement to the real cases to train segmentation models). We randomize the demographic features and clinical severity in GuideGen’s input while fixing the tumor location to a certain organ. It can be seen that on the multi-organ segmentation tasks, using 1000 synthetic cases as augmentation achieves an average performance gain of 6% across BTCV and AMOS datasets, while segmentation models trained using synthetic data alone achieves comparable performance to those trained with real data. For tumor segmentation tasks, performance gains are also observed using synthetic data as augmentation at a large scale, achieving a surplus in DSC of 9%, 1% and 5% separately for three different tumor types at 1000 synthetic cases. Using synthetic data for dataset synthesis also provides a solid starting point, especially for lung tumors, as seen in the last 6 rows of Tab. 14.
Dataset Method Std. DSC
Spl. Kid. Liver Sto. Pan. Adr. Eso. Aorta IVC Gall. Duo. Blad. PV&SV Avg.
BTCV Real 0.295 0.332 0.132 0.224 0.271 0.191 0.180 0.318 0.365 0.326 - - 0.208 0.258
5pt. MONAI-GM 0.013 0.018 0.011 0.090 0.098 0.070 0.124 0.073 0.077 0.310 - - 0.116 0.091
Zhuang’s 0.024 0.017 0.010 0.046 0.089 0.049 0.093 0.027 0.041 0.394 - - 0.156 0.086
MedDDPM 0.014 0.014 0.010 0.049 0.107 0.072 0.102 0.061 0.044 0.352 - - 0.141 0.088
MedSyn 0.015 0.012 0.007 0.027 0.113 0.086 0.111 0.062 0.031 0.398 - - 0.173 0.094
GuideGen 0.007 0.007 0.006 0.022 0.054 0.075 0.091 0.013 0.030 0.381 - - 0.062 0.068
AMOS Real 0.065 0.068 0.044 0.172 0.198 0.172 0.168 0.133 0.163 0.292 0.172 0.256 - 0.146
5pt. MONAI-GM 0.170 0.157 0.166 0.176 0.209 0.187 0.166 0.166 0.170 0.292 0.187 0.270 - 0.176
Zhuang’s 0.127 0.100 0.082 0.190 0.224 0.213 0.192 0.156 0.175 0.332 0.213 0.228 - 0.167
MedDDPM 0.062 0.043 0.062 0.199 0.192 0.204 0.186 0.090 0.123 0.304 0.204 0.231 - 0.141
MedSyn 0.070 0.045 0.087 0.103 0.172 0.149 0.163 0.085 0.131 0.281 0.149 0.252 - 0.127
GuideGen 0.025 0.041 0.017 0.132 0.147 0.152 0.141 0.084 0.088 0.271 0.152 0.240 - 0.112
Table 11: Segmentation variance on multi-organ segmentation tasks using real or synthetic cases from each method. The significance of our improvement over the second-best framework is tested using a two-sided $t$-test, with a $p$-value of 0.016 and 0.001 respectively for two datasets. Abbreviations: Spl.: Spleen, Kid.: Kidneys, Sto.: Stomach, Pan.: Pancreas, Adr.: Adrenal Glands, Eso.: Esophagus, IVC: Inferior Vena Cava, Gall.: Gallbladder, Duo.: Duodenum, Blad.: Urinary Bladder, PV&SV: Portal Vein and Splenic Vein. Some of the semantic classes reported are not consistent with the original dataset since they are merged as described in 6.1.2.
Method Std. DSC Std. $\text{HD}_{95}$
LU CO KI Avg. LU CO KI Avg.
Real 0.22 0.38 0.32 0.31 20.9 394.7 242.0 219.2
MONAI-GM 0.23 0.14 0.33 0.23 29.6 403.1 497.1 309.9
Zhuang’s 0.12 0.19 0.27 0.19 28.8 408.5 483.6 307.0
MedDDPM 0.21 0.19 0.38 0.26 21.2 409.0 275.9 235.4
MedSyn 0.35 0.18 0.29 0.27 45.5 411.4 426.0 294.3
GuideGen 0.21 0.10 0.40 0.24 20.7 357.2 465.7 281.2
Table 12: Segmentation variance on tumor segmentation tasks of different organs using real or synthetic cases from each method. The significance of our improvement over the second-best framework is tested using a two-sided $t$-test, with a $p$-value of 0.002, 0.01 and 0.06 respectively for three datasets. When there are no tumor predictions, HD95 is set to be 1000 to avoid $\infty$ values for a valid comparison. Abbreviations: LU: pulmonary cancer, CO: colorectal cancer, KI: renal cancer.
Figure 5: Samples generated by GuideGen for qualitative assessment. Five slices are shown in each case along with the multi-organ anatomies and tumor semantics (if any). We mark the bounding box of tumor labels in the CT slices shown using red outlines. Note that the first four cases accommodate a tumor site for the clinical description in the medical prompt while the last two has no tumor indication from the input prompt and generate a healthy anatomy and CT instead.
Figure 6: Qualitative results for multi-organ segmentations on BTCV and AMOS datasets. Three slices are shown in each case displayed.
Figure 7: Qualitative results for tumor segmentations on MSD-Lung, MSD-Colon and KiTS datasets. One slice is chosen for each case, which is centered at the largest transversal section of the tumor.
Classifier Type Metrics Classifier Classes
Accuracy AUROC
Age Group 0.74 0.69 [0,30], (30,40], (40,50], (50,60], (60,70], (70,80], (80$\infty$)
Race 0.76 0.64 Asian, Black or African American, White, Others
Gender 0.92 0.87 Female, Male
Tumor Location 0.89 0.73 Lung, Kidney, Colon and rectum
Table 13: Classifier performance on the validation split of TCIA and RJ datasets. Image-prompt alignment is assessed by evaluating the accuracy between features specified in input prompts and those predicted from generated images using these classifiers.
Dataset No. Real cases No. Synthetic cases for training / Avg. DSC
for training 0 100 200 500 1000
BTCV 0 0.00 0.69 0.79 0.79 0.81
24 0.74 0.81 0.82 0.83 0.83
AMOS 0 0.00 0.64 0.79 0.81 0.81
240 0.84 0.82 0.85 0.87 0.87
MSD-Lung 0 0.00 0.64 0.69 0.71 0.73
40 0.69 0.72 0.75 0.77 0.78
MSD-Colon 0 0.00 0.07 0.24 0.27 0.33
80 0.47 0.41 0.49 0.50 0.48
KiTS 0 0.00 0.46 0.51 0.59 0.63
210 0.72 0.72 0.72 0.74 0.77
Table 14: Performance gain by using different numbers of GuideGen’s generated samples for segmentation dataset synthesis or augmentation. We report the average DSCs for segmentation performance on each dataset. For reference, the main article reports the performance of segmentation models trained using all real training cases as opposed to those trained using 200 synthetic cases for each dataset.